OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = p + 1? (x, f(p)) (x, p + 1) (p + 1, f(p + 1)) (p + 1, f(p))

OpenStudy (anonymous):

as usual ... it is C the first coordinate is the input \(p+1\) and the second coordinate is the output \(f(p+1)\)

OpenStudy (haseeb96):

y=f(P+1) so as u know for graph we write (x,y) it show where the point it is so x=P+1 and y= f(P+1) now put it in (x,y) (P+1,f(P+1)) is the correct answer

jimthompson5910 (jim_thompson5910):

all ordered pairs are of the form (x,y) y is interchangeable with f(x), we now have (x,f(x)) finally, replace every copy of x with p+1 to get ( p+1, f(p+1) )

OpenStudy (anonymous):

Thank you all for the help :) I'd give you all medals if I could!